# references.bib

### v1

@@ 1 Introduction
The contributions of our work with BricksRL are threefold. First, we provide a unified platform integrating PyBricks and TorchRL within BricksRL, which facilitates the physical control of cost-effective robotic systems and the design of gym-like training environments for RL algorithms. This setup enables scalable training across a diverse array of robots, tasks, and algorithms. Second, we demonstrate how to extend the capabilities of BricksRL beyond the standard sensor set provided by LEGO. By integrating a camera sensor, we expand the platform’s capabilities, allowing for the creation of more diverse robots and tasks and thereby broadening its potential applications. Third, we present preliminary results that underscore the framework’s robustness and adaptability for real-world robotics applications. Furthermore, we provide explicit examples of sim2real transfer, demonstrating the practical utility of BricksRL in robotics and RL research. We make the source code available at: https://anonymous.4open.science/r/BricksRL. The building plans, and evaluation videos of the experiments are publicly available at https://bricksrl.github.io/ProjectPage/.
@@ 3.2 Environment
The PyBricksHubClass plays an important role in facilitating communication with the LEGO Hub, which controls the robots. It achieves this through the use of Bluetooth Low Energy (BLE) technology, which enables efficient, low-power communication. This class is designed to manage a two-way data exchange protocol, critical for meeting the real-time requirements of interactive applications. Importantly, PyBricksHubClass seamlessly bridges asynchronous communication with the traditionally synchronous structure of RL environments.
BaseEnv.
Given this basic structure, users can create their custom environments or extend existing environments for new tasks. All that needs to be done is to adapt the observation and action specifications and, if necessary, define a new reward function and adapt the step and reset function of the environment. In the appendix A.2.8 we share a simple template to create a custom environment.
Another benefit of using TorchRLs BaseEnv, apart from the simple integration with the RL agents of TorchRL, is the applicability of environment transforms, a fundamental feature of TorchRL. These transforms allow simple manipulation of the data received by the environment from the agent. TorchRL offers a broad spectrum of basic transforms such as frame stacking, observation normalization, and image transformations which can be especially useful for real-world robotics. Furthermore, the use of foundation models such as VIP [ma_vip_nodate] is made available through transforms with TorchRL and enhances the capabilities of experimentation with BricksRL. Detail descriptions of the environments that we implemented for the robots used in our experiments can be found in A.2.
@@ 3.3 LEGO Robots
Client Script.
All robots use the ByPricks Hubs for process management and control. Thereby each robot requires its own, specific client.py script in which the motors and sensors as well as the general workflow for processing and sending the data are defined. A new client script must be loaded and flushed to the Hub each time a new robot’s training begins, or when the Hub is replaced or utilized for a new robot, in case the underlying control flow changes or new sensors and motors are added or removed. For exact details provide an example of a typical client script that can be utilized as a template in the appendix 2.
@@
For effective control and communication between the robot and the environment, bidirectional data flow is essential. Therefore, we use the integration of Micropython in PyBricks on the robot side. Specifically, we make use of polling with standard input/output streams (stdin/stdout) to facilitate the seamless transmission of commands and sensory feedback between the robot and the environment.
Actions determined by the agent are encoded into a byte-stream format and transmitted to the robotic system. On the robot, the streamed data is getting polled with a registered stdin stream. This monitoring system enables the robot to detect and parse incoming command signals efficiently. Once parsed, signals are applied to the robot’s motors. Concurrently, the sensory modules collect data on its current state and environmental interactions, which are then encoded into a byte stream formatted for transmission back to the environment via stdout.
Similarly, PyBricks employs a polling mechanism, established by registering the stdin stream for incoming data. This setup ensures that the class is constantly aware of new commands being transmitted from the robot. The incoming data is queued and sequentially processed to extract the sensory data of the robot. Once the data has been extracted and converted into floating points, it is added to the TensorDict and made available to the agent to evaluate the current state and select a new action 1.
@@ 5 Conclusion
In this paper, we introduced BricksRL and detailed its benefits for robotics, RL, and educational applications, emphasizing its cost-effectiveness, reusability, and accessibility. Additionally, we showcased its practical utility through the deployment of three distinct robots, performing various tasks with a range of sensors, across more than 100 experiments. Our results underscore the viability of integrating state-of-the-art RL methodologies within both research and educational contexts via BricksRL. By providing comprehensive building plans and facilitating access to BricksRL, we aim to establish this investigation as a foundational proof of concept for utilizing LEGO-based robots in the training of RL algorithms. Looking ahead, promising avenues for exploration include the creation of more complex robots and tasks as well as the introduction of datasets for imitation learning and Offline-RL.
@@ A.2.4 WalkerSim-v0
Action and observation specifications are the same as in the Walker-v0 6.
@@ A.2.6 RoboArmSim-v0
Action and observation specifications are the same as in the RoboArm-v0 7.
@@ A.3 Client Script
For each task and robot, a custom client script is required to facilitate the interaction between the robot and the environment. Listing 2 provides a straightforward example of such a script tailored for the RunAway-v0 task. In this example, a single float value representing the action is processed to control the motors. Subsequently, sensor data are collected and transmitted back to the environment.

### v2

@@ 1 Introduction
The contributions of our work with BricksRL are threefold. First, we provide a unified platform integrating Pybricks and TorchRL within BricksRL, which facilitates the physical control of cost-effective robotic systems and the design of gym-like training environments for RL algorithms. This setup enables scalable training across a diverse array of robots, tasks, and algorithms. Second, we demonstrate how to extend the capabilities of BricksRL beyond the standard sensor set provided by LEGO. By integrating a camera sensor, we expand the platform’s capabilities, allowing for the creation of more diverse robots and tasks and thereby broadening its potential applications. Third, we present preliminary results that underscore the framework’s robustness and adaptability for real-world robotics applications. Furthermore, we provide explicit examples of sim2real transfer and the application of datasets with offline RL, demonstrating the practical utility of BricksRL in robotics and RL research. We make the source code available at: https://github.com/BricksRL/bricksrl. The building plans, and evaluation videos of the experiments are publicly available at https://bricksrl.github.io/ProjectPage/.
@@ 3.2 Environment
Developed by BricksRL, the PybricksHub class plays an important role in facilitating communication with the LEGO Hub, which controls the robots. It achieves this through Bluetooth Low Energy (BLE) technology, which enables efficient, low-power communication. This class is designed to manage a two-way data exchange protocol, critical for meeting the real-time requirements of interactive applications. Importantly, the PybricksHub class seamlessly bridges asynchronous communication with the traditionally synchronous structure of RL environments.
EnvBase.
In BricksRL, environments are designed as classes that inherit from EnvBase provided by TorchRL, which is a foundational class for building RL environments. This structure gives access to the TorchRL ecosystem and simplifies the creation of environments. Users can create custom environments or extend existing environments for new tasks. All that needs to be done is to adapt the observation and action specifications and, if necessary, define a new reward function and adapt the step and reset function of the environment.
A key advantage of using TorchRL’s EnvBase in BricksRL is the ability to apply environment transforms, a fundamental feature of TorchRL. These transforms enable simple manipulation of the data exchanged between the environment and the agent. TorchRL provides a wide range of useful transforms, such as frame stacking, observation normalization, and image transformations, which are particularly valuable for real-world robotics. Additionally, the integration of foundation models like VIP [ma_vip_nodate] through transforms expands the experimentation capabilities within BricksRL. Detailed descriptions of the environments we implemented for our experiments, along with a template for creating custom environments, can be found in A.2 and A.2.8, respectively.
@@ 3.3 LEGO Robots
BricksRL employs a bidirectional data flow between the robot and the environment, facilitated by MicroPython in Pybricks. The agent’s actions are transmitted as byte streams via standard input (stdin), where they are parsed and applied to the robot’s motors. At the same time, the robot sensor data is sent back to the environment through standard output (stdout) for state evaluation and action selection. Each robot uses a dedicated client script to manage its motors, sensors, and control flow for specific tasks. For exact details and an example of a typical client script, see the provided template in A.3.
@@ 4.3 RoboArm
4.4 Offline Training
Offline RL uses pre-collected datasets to train algorithms efficiently, avoiding real-world interactions and complex simulations. To further highlight the capabilities of BricksRL for education and research in robotics and RL we collected offline datasets for the LEGO robots. With those datasets, BricksRL allows training of the LEGO robots via offline RL or imitation learning, both of which are state-of-the-art methods for RL in robotics.
For BricksRL, we curated datasets for three robot configurations: 2Wheeler, Walker, and RoboArm. These datasets include both expert and random data for four tasks in our experiments (Walker-v0, RoboArm-v0, RunAway-v0, Spinning-v0). Details about the datasets and dataset generation can be found in the appendix A.7. Using these datasets, we demonstrated that offline RL with BricksRL is feasible, successfully training both online and offline RL algorithms and applying them to a real robot. The evaluation performance, shown in Table 4, highlights the superior performance of offline RL algorithms, particularly with expert data, while online algorithms struggle, suggesting overfitting or poor generalization. For further details on training parameters, please refer to the appendix A.8.
Walker-v0 RoboArm-v0 RunAway-v0 Spinning-v0
Agent Random Expert Random Expert Random Expert Random Expert
TD3 $-79.71$ $-153.45$ $-124.22$ $-201.94$ 19.86 $9.06$ 6160.25 6168.28
SAC -- 66.91 $-255.41$ $-54.67$ $-218.76$ $14.74$ $10.80$ 5416.11 9349.52
BC $-202.46$ $-85.65$ $-117.72$ $-7.34$ $-0.27$ 18.13 35.55 9150.02
IQL $-136.13$-74.80 $-76.89$-3.10 14.07 18.80 4544.28 9096.60
CQL $-72.75$ $-77.93$-46.91 $-17.41$ 19.60 19.74 4509.31 9099.03
Table 4: Evaluation Results: Online (TD3, SAC) and Offline (BC, IQL, CQL) RL Algorithms. Scores represent the mean reward averaged over 5 episodes and 5 random seeds.
@@ 5 Conclusion
In this paper, we introduce BricksRL and detail its benefits for robotics, RL, and educational applications, emphasizing its cost-effectiveness, reusability, and accessibility. In addition, we showcased its practical utility by deploying three distinct robots, performing various tasks with a range of sensors, across more than 100 experiments. Our results underscore the viability of integrating state-of-the-art RL methodologies through BricksRL within research and educational contexts. By providing comprehensive building plans and facilitating access to BricksRL, we aim to establish this investigation as a foundational proof of concept for utilizing LEGO-based robots to train RL algorithms. Moving forward, avenues for further research include creating more complex robots and tasks, exploring applications in multi-agent settings, and leveraging large datasets to enhance RL training through transformer-based imitation learning. Ultimately, BricksRL sets the stage for a future where accessible, reusable robotic systems support and expand RL research, collaborative learning, and interactive education.
Acknowledgements
We thank A. De Fabritiis Campos, M. De Fabritiis Campos and P. Vallecillos Cusco for providing their LEGOs to this project.
@@ A.2.4 WalkerSim-v0
Action and observation specifications are the same as in the Walker-v0 7.
@@ A.2.6 RoboArmSim-v0
Action and observation specifications are the same as in the RoboArm-v0 8.
@@ A.3 Client Script
For each task and robot, a custom client script is required to facilitate interaction between the robot and the environment. The client.py script defines the configuration of motors, sensors, and the workflow for processing and exchanging data. This script must be uploaded to the Pybricks Hub and updated whenever the robot’s configuration changes, such as when motors or sensors are added or removed. Listing 2 provides a simple example of a client script tailored for the RunAway-v0 task. In this example, a single float value representing the action is used to control the motors, while sensor data is collected and transmitted back to the environment.
@@ A.6 Online Training Parameter
A.7 Dataset Generation Process
The expert dataset for each robot configuration was generated by training a Soft Actor-Critic (SAC) agent to solve the respective task and record transitions over 100 episodes on the real robot. The random dataset was created by executing a random policy for 100 episodes. For example, the collection process for the Walker robot took about an hour, yielding approximately 10,000 transitions. Details such as mean reward, number of transitions, and collection episodes for each dataset can be found in Table 11. The dataset is available on Hugging face.
Task Mean Reward Expert Transitions Random Transitions Episodes
Walker-v0 $-69.12$ $9,244$ $10,000$ $100$
RoboArm-v0 $-9.87$ $1,297$ $10,000$ $100$
RunAway-v0 $8.041$,98719871,9871 , 987 $1,612$ $100$
Spinning-v0 $8981.19$ $5,000$ $5,000$ $100$
Table 11: Dataset Statistics
A.8 Offline Training Parameter
For the online algorithms, we used the same parameters in our offline rl experiments as in the online experiments.
Parameter BC IQL CQL
Learning Rate (lr) $3\times 10^{-4}$ $3\times 10^{-4}$ $3\times 10^{-4}$
Batch Size 256 256 256
Number of Cells 256 256 256
Gamma - 0.99 0.99
Soft Update $\epsilon$ 0.995 0.995
Loss Function L2 L2 L2
Temperature - 1.0 1.0
Expectile - 0.5 -
Min Q Weight - - 1.0
Max Q Backup - - False
Deterministic Backup - - False
Num Random Actions - - 10
With Lagrange - - True
Lagrange Threshold - - 5.0
Normalization LayerNorm None None
Dropout 0.01 0.0 0.0
BC Steps - - 1,000
Table 12: Hyperparameter for the agents BC, IQL, and CQL
We trained the models for various tasks and datasets with different update counts 13. The RunAway-v0 task was trained for 2,000 updates on both the expert and random datasets. For the Spinning-v0 task, we used 5,000 updates across both datasets. The Walker-v0 task required 10,000 updates for both expert and random datasets. Finally, the RoboArm-v0 task was trained for 10,000 updates on the random dataset and 5,000 updates on the expert dataset.
Task Expert Random
RunAway-v0 2,000 2,000
Spinning-v0 5,000 5,000
Walker-v0 10,000 10,000
RoboArm-v0 5,000 10,000
Table 13: Number of offline training updates for each task and dataset
A.9 Network Architecutre
Throughout the experiments, all algorithms utilize the same architecture for the policy, Q-functions, and value functions (where applicable). Each network is structured as a three-layer multilayer perceptron (MLP), with specific TorchRL actor modules used for the policy, depending on the algorithm. For more details, we refer readers to the code repository: GitHub.
The only variation in architecture occurs when incorporating pixel-based observations. In this case, a convolutional neural network (CNN) is used to encode the image data. These encodings are then concatenated with the sensor-based encodings, and the combined embeddings are passed through a shared MLP. Detailed implementation specifics can be found in the GitHub repository.